OpenStudy (fairygirl0404):

Which point is NOT on the graph h(x)= ∛(x+64) (-64, 0) (-72, -2) (65, 5) (-37, 3)

OpenStudy (fortytherapper):

Have you attempted to plug in the answer choice values?

OpenStudy (fortytherapper):

Also know that \[\sqrt[n]({Expression})^m = (Expression)^\frac{ m }{ n }\]

OpenStudy (mathmale):

@fairygirl0404: Would you kindlyl acknowledge the help others have given you, and let them know when you consider the posting closed.

OpenStudy (fortytherapper):

She closed it? She's not my fairy I guess =(

OpenStudy (fairygirl0404):

@FortyTheRapper...I'm not sure what happened I did not close my question. And I did what you suggested and got my answer (65,5). Thank you for your help. @mathmale I have acknowledge him.

OpenStudy (fortytherapper):

@fairygirl0404 Sounds good, nice job!